OpenStudy (anonymous):

for a circle with a radius of 8 ft find the arc lengths subtented by a central angle of pi/3

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {x \over 360} \times Circumference }\) \(x\) is the angle, which is pi/3 here.

Parth (parthkohli):

I suggest converting it to degrees first, or doing it with radians only.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {x \over 2\pi} \times Circumference }\)

OpenStudy (anonymous):

okay thanks! :)